16.2 Calculate the pH and percent dissociation for the following reactions of acids or bases with water.

b. 50 mmol of acetic acid plus 20mmol of sodium acetate in 250 mL of water
Answer #1

acetic acid an sodium acetate forms a buffer

according to Handerson Hasselbach equation

pH = pKa + log [salt] / [acid]

pKa of acetic acid = 4.76

pH = 4.76 + log [20] / [50] = 4.36

hence [H+] = 10^-pH = 10-4.36 =0.000044